### v2

# NESTful: A Benchmark for Evaluating LLMs on Nested Sequences of API Calls

Abstract
The resurgence of autonomous agents built using large language models (LLMs) to solve complex real-world tasks has brought increased focus on LLMs’ fundamental ability of tool or function calling. At the core of these agents, an LLM must plan, execute, and respond using external tools, APIs, and custom functions. Research on tool calling has gathered momentum, but evaluation benchmarks and datasets representing the complexity of the tasks have lagged behind. In this work, we focus on one such complexity, nested sequencing, with the goal of extending existing benchmarks and evaluation. Specifically, we present NESTful, a benchmark to evaluate LLMs on nested sequences of API calls, i.e., sequences where the output of one API call is passed as input to a subsequent call. NESTful contains 1800+ nested sequences where all the function calls are executable. Experimental results on multiple models and settings show that the best-performing model on the dataset has a full sequence match accuracy of 25% and win-rate of 34% necessitating a large scope for improvement in the nested sequencing aspect of function calling. Our analysis of these results provides possible future research directions for the community, in addition to a benchmark to track progress. We have released the NESTful dataset under the Apache 2.0 license at https://github.com/IBM/NESTFUL.
NESTful: A Benchmark for Evaluating LLMs
on Nested Sequences of API Calls
Kinjal Basu1∗, Ibrahim Abdelaziz1∗, Kiran Kate1, Mayank Agarwal1, Maxwell Crouse1, Yara Rizk1, Kelsey Bradford2, Asim Munawar1, Sadhana Kumaravel1, Saurabh Goyal1, Xin Wang1, Luis A. Lastras1, and Pavan Kapanipathi1 1IBM Research, USA           2Georgia Institute of Technology {kinjal.basu, ibrahim.abdelaziz1, mayank.agarwal, maxwell.crouse, yara.rizk}@ibm.com {asim, sadhana.kumaravel1, saurabh.goyal1, chloe.wang }@ibm.com kelseymaria5@gmail.com, {kakate, lastrasl, kapanipa}@us.ibm.com
**footnotetext: These authors contributed equally to this work
1 Introduction
Autonomous agents, built with Large language models (LLMs), are gaining popularity in solving complex, real-world problems Yao et al. (2023); Deng et al. (2024). LLMs handle a user’s request by understanding their intents, planning the required tasks to address it, executing those tasks step by step, and providing a response. For most real-world problems Jimenez et al. ; Roy et al. (2024); Thakur et al. (2023), LLMs must interact with external environments through tool, function, and API calls (Application Programming Interface), which primarily leverages LLMs’ tool calling abilities [^1] API, function calling and tool-use are used interchangeably throughout the paper.
The significant reliance on LLMs’ function calling abilities led recent research to continuously improve this dimension of LLMs. On one hand, approaches to improve function calling have exploded Abdelaziz et al. (2024); Liu et al. (2024); Srinivasan et al. (2023); on the other hand, benchmarks and evaluations are lagging behind. For instance, BFCL v1 and v2 focused on evaluating single, multiple, and parallel function calling tasks for both non-executable and executable versions Patil et al. (2023). Works such as API-Blend Basu et al. (2024) complemented prior work by introducing granular task evaluation such as slot-filling, API-detection, and sequencing. BFCL v3 has progressed further into agentic use cases with multi-step and multi-turn function calling evaluation.
However, benchmarks for fundamental but complex tasks such as the sequencing of functions have not been well explored yet, which forms the basis of this work. Existing evaluation benchmarks pose sequencing as the prediction of single or multiple isolated API calls, where the output of any particular API call within that sequence is considered irrelevant. In contrast, for many real-world tasks, a sequence of API calls is nested, i.e., the output of some API calls is used in the arguments of subsequent API calls. Figure 1 shows one such example of a nested sequence of APIs.
Figure 1: End-to-End Evaluation Pipeline: NESTful provides a test set of input queries and its corresponding list of nested function calls. It also provides executable implementations for each tool in the library and allows for evaluating models in direct prompting or REACT styles. Given a query, the pipeline infers the input LLM to generate the required sequence of function calls, execute those functions (taking into account nested variables) and compare the final answer with the gold one.
In this paper, we present NESTful, a benchmark specifically designed to evaluate models on nested API calls and it contains over 1800 nested sequences. The data consists of: (1) user queries, (2) a catalog of APIs and their specifications, (3) the sequence of API calls and the corresponding parameters, and (4) the expected output response. The datasets are based on the MathQA Amini et al. (2019) and StarCoder2-Instruct Wei et al. (2024) datasets which are commonly used in the literature but are missing the executable component. Figure 1 shows the end-to-end evaluation pipeline of NESTful.
We evaluated the dataset on 15 state-of-the-art models from the literature and exposed the gaps of these models in handling complex function calling sequences. The best model did not exceed 25% full sequence match nor 33% on win rate metrics. Models struggled as the nesting got deeper and the data dependencies increased. To further advance research in this area, we will publicly release the NESTful dataset with executable Python implementation for each tool and the evaluation code for the models.
2 Related Work
The best way to enable API function calling in LLMs remains an active area of research. Methods that utilize large, general-purpose proprietary models (e.g., Gemini (Team et al., 2023) or GPT (Achiam et al., 2023)) typically make use of carefully constructed prompts and in-context learning examples, e.g., Song et al. (2023). Smaller, more specialized models often start from a strong-performing code model (e.g., DeepSeek-Coder (Guo et al., 2024), CodeLlama (Roziere et al., 2023), or Granite Code (Mishra et al., 2024)) and fine-tune primarily on highly curated datasets Srinivasan et al. (2023); Ji et al. (2024); Abdelaziz et al. (2024) that have been extended with synthetic data Zhang et al. (2024a).
Most of these existing works initially focused on basic function calling abilities that did not involve much complexity. Recent advances in enabling LLMs to handle complex multi-API interactions have introduced structured methods like Reverse Chain Zhang et al. (2024b). This approach employs backward reasoning to optimize multi-step API planning, allowing LLMs to effectively manage nested workflows by aligning intermediate steps with the final goal. Such methods highlight LLMs’ potential to perform efficient, target-driven planning in resource-constrained environments.
To evaluate and enhance these aforementioned approaches, numerous works released training and benchmarking data in service of API function calling, such as ToolLLM Qin et al. (2023); APIBench Patil et al. (2023); APIGen Liu et al. (2024); or API-BLEND (Basu et al., 2024). While those mostly focused on simpler tasks (e.g., single or multiple isolated API calling functionality), NesTools (Han et al., 2024) focused on more complex real-world applications. It was designed to evaluate the nested tool learning capabilities of LLMs, addressing the lack of nested sequencing related data in the existing benchmarks. Similarly, our work focuses on the subset of function calling tasks that need an interdependent sequence of API calls, a necessity in many real-world, multi-step problems.
Unlike NesTools which was created through a fully synthetic process, NESTful is derived from well-known existing datasets. Additionally, the average sequence length for NESTful is 4.36, compared to 3.04 for NesTools. Other function-calling benchmarks, such as SealTool Wu et al. (2024) and BFCL-v3222https://gorilla.cs.berkeley.edu/blogs/13_bfcl_v3_multi_turn.html, include some nested sequences but are smaller than NESTful as they are not specifically designed for nested function calls.
Finally, AgentBoard (Ma et al., 2024) is a comprehensive benchmark and open-source evaluation framework designed for assessing LLMs as general-purpose agents. It introduces fine-grained progress metrics and interactive visualization tools to analyze incremental advancements and multi-faceted agent behaviors, addressing the limitations of traditional success-rate-focused evaluations. This framework enhances interpretability and fosters the development of more robust LLM agents. In contrast, our NESTful benchmark is more focused on the evaluation of tool-augmented LLMs and calculates granular offline metrics (F1 score, accuracy, etc.) in addition to the win rate.
3 NESTful Dataset Curation
Figure 2: NESTful Data Schema for Input, Tool Library, Output, and Python implementation. In Output JSON, the arg_2 of Function_Name_2 showcases how the variable assignments are used to create a nested sequence of function calls.
NESTful comprises of more than 1,800 instances designed for benchmarking tool-augmented LLMs, with a focus on nested sequencing. Each instance consists of (1) a user query, (2) a list of all available tools for the model to choose from, (3) the gold sequence of tools and their arguments needed to answer the user query, and (4) the final answer that should be obtained once the tools are executed. The dataset also contains corresponding Python code for every API in the library as well as a mechanism to run the input query via any LLM, execute the tools predicted by the LLM, and provide the final answer.
3.1 Nested Function Calling Data Schema
The NESTful data schema demonstrated in Figure 2 showcases the template used for representing the Input, Tool Library, Output, and Python Code. An important aspect of nested function calling is to enable a mechanism for tool reference; i.e. a subsequent tool call using that reference to access the output of the previous tool execution. To do so, we assign a unique variable name to each tool which distinctly identifies each tool, even when two identical tools with different arguments appear in the same sequence (parallel API calls). For example, in Figure  1, “rectangle_area” tool was assigned “label”: “\$var_1”. This allows the next tool “multiply” to use the output of “rectangle_area” as an argument: “arg_0”: “\$var_1.result\$”.
3.2 NESTful Data Domains
NESTful is composed of data from two domains; 1) mathematical reasoning data and 2) generic tools from the coding domain. We describe below the process followed to create each data category.
3.2.1 Mathematical Reasoning Data
For the first part of NESTful, we relied on datasets that test the model for nested function calling in the math domain. In particular, we build on MathQA Amini et al. (2019); a benchmark designed for evaluating the ability of AI models to solve mathematical word problems. It consists of questions that test numerical reasoning and problem-solving skills, requiring models to both understand the text of a word problem and perform mathematical operations to arrive at the correct solution. An example is shown in Figure 5.
Tool specifications
Since MathQA provides only the tool names, we manually created specifications for all the tools in the dataset. This covers 40 tools in total; e.g. divide, remainder, volume_cylinder, permutation, etc. For each tool, we define the name, tool description, and detailed outline of the tool input and output parameters including the parameter data type and description as shown in Figure 3.
Figure 3: Sample specification for tools derived from MathQA
Tool calling input-output pairs
To build the test data, we used the test set of MathQA where the “problem” definition is the query and parsed the “annotated formula” into a nested sequence of tool calls. An example is shown in Figure 5.
Figure 4: [top] An example function (angle_rotation) from StarCoder2-Instruct dataset with its docstring documentation. [bottom] Tool specification for “angle_rotation” after inferring the different data types and creating its input/output parameters.
Figure 5: Sample problem from the MathQA dataset. The Operation Stack provides an ordered sequence of nested tool calls which we transform into a Tool Call format for the NESTful dataset.
Executable Code and Filtering
For each tool, we also generated its corresponding implementation in Python. This allows us to execute the nested call sequence and match the execution result with the gold answer. This ensures the correctness of the set of corresponding tools and the code execution too. We then filtered out any samples where we could not reproduce the gold answer from executing the nested tool calls. This process resulted in 1,415 test samples covering 40 tools with an average 5.1 tool calls per sample.
3.2.2 Coding Data
We also curated test examples based on generic Python functions from the StarCoder2-Instruct dataset Wei et al. (2024). This dataset has a total of 50K Python functions and covers a wide range of tools that can be used.
Tool specifications
We leveraged StarCoder2-Instruct’s Python implementations and docstrings (such as in the example shown in Figure 4) to create API specifications. For each Python function, (1) we used an LLM (Mixtral 8x22B [^3] https://huggingface.co/mistralai/Mixtral-8x22B-Instruct-v0.1) to generate the possible Python types of the input and output arguments, and then (2) we validated and executed the function code itself to ensure it does not produce any errors. After both steps, we generated a corresponding JSON specification for each function documenting its input-output arguments. Figure 4 shows an example for “angle_rotation” function.
Tool calling input-output pairs
To create input-output pairs, we synthetically generated 446 nested tool calling instances. To do so, we leveraged DiGiT [^4] https://github.com/foundation-model-stack/fms-dgt synthetic data generation framework where we manually curated 10 seed examples and used Mixtral-8x22B as the teacher model to generate the rest. We also built a validator to check the correctness of the produced samples w.r.t. ensuring the tools and parameters used are not hallucinated, required parameters are specified and there is at least one nested tool call in the output sequence.
Filtering
We further filtered the generated input-output pairs by executing their gold nested API sequence to ensure they execute and attach the result as the gold answer. For a function that has a randomness element (e.g., generating a random list), we set a fixed seed for all our experiments and re-execute all those cases to ensure that we are getting the same response all the time. From this category of data, we generated 446 test examples covering more than 881 distinct tools with an average tool sequence length of 2.1.
4 Evaluation
Model Total Parameters One-shot ICL Three-shot ICL
F1 Func. F1 Param. Part. Acc. Full Acc. Win Rate F1 Func. F1 Param. Part. Acc. Full Acc. Win Rate
xLAM-1b-fc-r 1B 0.35 0.16 0.13 0.04 0.03 0.49 0.20 0.15 0.06 0.04
xLAM-7b-fc-r 7B 0.54 0.32 0.18 0.10 0.13 0.49 0.35 0.23 0.18 0.19
MadeAgents/Hammer2.0-7b 7B 0.62 0.42 0.29 0.25 0.33 0.61 0.46 0.31 0.25 0.34
granite-3.0-8b-instruct 8B 0.50 0.27 0.15 0.05 0.07 0.52 0.35 0.22 0.15 0.16
llama-3-1-8b-instruct 8B 0.58 0.37 0.24 0.18 0.16 0.53 0.38 0.24 0.19 0.16
Team-ACE/ToolACE-8B 8B 0.51 0.23 0.13 0.00 0.00 0.51 0.27 0.14 0.00 0.00
llama-3-2-11b-instruct 11B 0.57 0.42 0.26 0.21 0.18 0.60 0.38 0.26 0.21 0.17
Granite-20B-FunctionCalling 20B 0.60 0.37 0.28 0.24 0.26 0.59 0.35 0.26 0.22 0.24
Mixtral-8x7B-Instruct-v0.1 46.7B 0.29 0.20 0.14 0.12 0.11 0.36 0.25 0.16 0.13 0.13
xLAM-8x7b-fc-r 46.7B 0.44 0.27 0.16 0.01 0.01 0.49 0.29 0.17 0.02 0.03
llama-3-1-70b-instruct 70B 0.35 0.18 0.10 0.00 0.01 0.33 0.16 0.10 0.00 0.03
llama-3-2-90b-instruct 90B 0.32 0.16 0.09 0.00 0.02 0.29 0.14 0.10 0.00 0.03
Mixtral-8x22B-Instruct-v0.1 141B 0.43 0.30 0.22 0.16 0.10 0.64 0.48 0.29 0.21 0.29
xLAM-8x22b-fc-r 141B 0.59 0.40 0.23 0.15 0.08 0.59 0.40 0.23 0.14 0.11
llama-3-1-405b-instruct-fp8 405B 0.21 0.20 0.13 0.10 0.08 0.26 0.25 0.15 0.11 0.14
Table 1: Evaluation results on NESTful on state-of-the-art LLMs with Direct Prompting technique. Models are sorted by their size. Experiments are done in one-shot and three-shot ICL settings. The best performance is highlighted in bold; the second best is underlined. Partial Sequence Accuracy (Part. Acc.) denotes the percentage of calling the correct API sequence (API names and arguments), whereas Full Sequence Accuracy (Full Acc.) counts the percentage of times where the model gets the entire sequence of APIs correctly. Both scores range from 0 to 1. We also report Win Rate which measures whether all the predicted APIs by the model are executable and lead to an exact match with the gold answer.
4.1 Baselines
We evaluated NESTful using 15 open-source models with sizes varying from 1B up to 405B parameters. This list includes the top models on Berkeley Function-Calling Leaderboard (BFCL) [^5] https://gorilla.cs.berkeley.edu/leaderboard.html. xLAM-1b-fc-r and xLAM-7b-fc-r (Liu et al., 2024) are fine-tuned function calling models based on DeepSeek-Coder-1.3B-instruct and DeepSeek-Coder-7B-instruct-v1.5. Similarly, xLAM model family also include xLAM-8x7b-fc-r [^6] https://huggingface.co/Salesforce/xLAM-8x7b-r and xLAM-8x22b-fc-r [^7] https://huggingface.co/Salesforce/xLAM-8x22b-r based on Mixtral 8x7B and Mixtral 8x22B, respectively. We also evaluate other competitive recent function calling models like Hammer2.0-7b [^8] https://huggingface.co/MadeAgents/Hammer2.0-7b and ToolACE-8B [^9] https://huggingface.co/Team-ACE/ToolACE-8B. We compare against various sizes of the latest LLAMA 3.1 models; particularly, llama-3-1-8b-instruct, llama-3-2-11b-instruct, llama-3-1-70b-instruct, llama-3-2-90b-instruct, llama-3-1-405b-instruct Dubey et al. (2024). We finally evaluate against other open models such as Granite-3.0-8b-instruct Granite Team (2024) and their tool calling variant Granite-20B-FunctionCalling Abdelaziz et al. (2024).
4.2 Experimental Settings
The experiments are carried out in one-shot and three-shot settings, i.e., the prompt contains one or three in-context learning (ICL) examples, respectively. To the best of our knowledge, all the 15 models were not trained with the label assignment syntax in the output API sequence, so it was crucial to have ICL examples to get the best results from the models. For each model, we used its specified prompt along with the special tags. Context length limitations prevented the inclusion of the entire API library in the prompt. Instead, we pre-processed the data to create a shorter API list for each sample. This list ensured the inclusion of the gold APIs, the APIs used in the in-context learning (ICL) examples, and some random APIs, keeping the total prompt length under 4k tokens (based on granite-3.0-8b-instruct tokenizer). Also, the API calls were extracted from the model’s response as a list of JSON objects, taking into account the specific prompt format and output structure for each model. Finally, we evaluated a zero-shot ReAct Yao et al. (2022) agent with the best 4 models. The Think-Act-Observe loop in ReAct was limited to 10 iterations. All experiments have employed greedy sampling and set the temperature to its lowest value of 0.0 for all the models.
4.3 Metrics
For a detailed evaluation, we use the following metrics: 1) F1 score for function and parameter names generation Abdelaziz et al. (2024), 2) Partial and Full Match Accuracy, and 3) Win Rate.
Generated response from the model is a sequence of API calls, with each call consisting of an API name and its argument-value pairs. We use the Partial Sequence Matching metric to determine how many predicted APIs (with their argument-value pairs) in a sequence match with the gold API sequence. In contrast, the Full Sequence Matching metric evaluates whether the model predicts the exact full sequence of APIs, including both the API names and their argument-value pairs, when compared to the gold API sequence. In both cases, we calculate the scores for each sample and then compute the statistical mean across the entire dataset as the final score. The final metric we use is the Win Rate, which measures if all the predicted APIs by the model are valid and when executed lead to the gold answer. Unlike the F1 and Accuracy measures, which assess the alignment of predictions with the gold API sequence, the Win Rate focuses on the final gold answer. A win is recorded if the predicted answer matches the gold answer.
Model Total Parameters Direct Prompting (One-shot ICL) ReAct Agent (Zero-shot)
F1 Func. F1 Param. Part. Acc. Full Acc. Win Rate F1 Func. F1 Param. Part. Acc. Full Acc. Win Rate
MadeAgents/Hammer2.0-7b 7B 0.62 0.42 0.29 0.25 0.33 0.56 0.34 0.14 0.00 0.25
Granite-20B-FunctionCalling 20B 0.60 0.37 0.28 0.24 0.26 0.40 0.00 0.00 0.00 0.11
llama-3-1-70b-instruct 70B 0.35 0.18 0.10 0.00 0.01 0.30 0.36 0.13 0.00 0.42
Mixtral-8x22B-Instruct-v0.1 141B 0.43 0.30 0.22 0.16 0.10 0.31 0.38 0.17 0.02 0.52
Table 2: Evaluation results on NESTful comparing the performance of a ReAct Agent (zero-shot) to the Direct Prompting with a one-shot ICL example. The best performance is highlighted in bold; the second best is underlined.
4.4 Results
Table 1 presents a comparison of different baselines on the NESTful dataset with one-shot and three-shot ICL example settings. The low numbers of the best function calling models in Table 1 depicts the complexity and toughness of the nested sequencing problem. Hammer2.0-7b achieved the highest performance with a full accuracy of 25% and a win rate of 34%, significantly below acceptable numbers for real world applications. These numbers are similar to multi-turn setting in BFCL v3 dataset. [^10] https://gorilla.cs.berkeley.edu/leaderboard.html This clearly depicts the significant scope for improvements for the models in various aspects of function calling including nested sequencing. We inspected the models’ outputs and identified several common issues across them. These models struggle with tasks such as assigning variables, utilizing output parameter details from the API specifications, and correctly passing variable names and corresponding output parameters to subsequent APIs, even with ICL examples [^11] Note: While we acknowledge that the models aren’t trained with the robust data schema discussed in Section 3.1, common issues with nested sequencing immaterial of the data schema has to improve in LLMs..
As anticipated, in most of our experiments in Table 1, the models are doing better across all the metrics when they are provided with three ICL examples in the prompt instead of one example. Across all models, Partial Sequence Match scores are consistently higher than and Full Sequence Match scores, which is expected, as the latter is a stricter metric than the former. In some cases, the Win Rate is a little higher than the Full Accuracy because models sometimes follow alternate trajectories to reach the final answer. Figure 6 shows this case where a correct final answer can be obtained by following two different solution paths. Despite the smaller model sizes, the function calling models such as Hammer2.0-7b, Granite-20B-FunctionCalling, and xLAM series did exceptionally well compared to the larger language models. Across all the metrics and evaluation settings (one shot and three shots), Hammer2.0-7B achieved either the highest or the second-highest scores. On the contrary, being a function calling model, ToolACE-8B still suffered in following the output JSON format with the proper variable assignments and that led to very bad accuracy and win-rate scores. For this benchmark, we found that the size or complexity of the model itself becomes less relevant, as long as it can effectively follow the given instructions and the ICL examples. As evidence, we can see that some larger models such as xLAM-8x22b-fc-r and llama-3-1-405b-instruct-fp8 performed poorly, whereas a 7B model (Hammer2.0-7b) was doing best.
Figure 6: An example where the predicted API sequence is valid but does not match the ground truth sequence. The final answer match results in a Win Rate of 1.
ReAct-based Evaluation:
The previous results showed how poorly direct prompting of LLMs performed on NESTful. The literature has shown that agentic approaches have resulted in better performance on complex tasks. While many agentic architectures exist, we selected ReAct due to its popularity and its ability to reason over the output of the tool that is added to the prompt at each turn.
Table 2 summarizes the results of the ReAct agent Yao et al. (2023) in comparison to the one-shot ICL direct prompting setting. Note that there was no ICL example provided in the ReAct case as the expected output does not need to follow the label assignment syntax. The reported metrics (except Win Rate) which measure the conformance to the ground truth API sequence show that ReAct performed poorly. This could be due to the single-step planning and execution approach as opposed to planning the entire API sequence at once.
ReAct does outperform direct prompting on Win Rate but has room for improvement. Inspecting the outputs showed that the top two models had better format alignment to ReAct and sometimes performed the functionality of some of the APIs using its parametric knowledge. This led to lower F1 and accuracy scores but higher Win Rate; the computations were correct despite mismatching steps, as shown in Figure 7.
Figure 7: An example where a ReAct agent using llama-3-1-70b-instruct used its parametric knowledge to perform a computation correctly as highlighted. This led to a final answer that matches the ground truth answer but the API sequence does not match with the ground truth sequence.
4.5 Dataset Analysis
To analyze the results, we model the samples in the NESTful as a Directed Acyclic Graph (DAG) where nodes are individual function calls and the edges are data dependencies between two nodes.
(a) Data Nesting Depth Analysis
(b) Total Data Dependencies Analysis
(c) Nesting Pattern Analysis
Figure 8: Model performances with a varying number of (a) longest data dependencies, (b) the total number of data dependencies, and (c) common data nesting patterns in NESTful. We observe that models perform well for the simple data samples but the performance drops sharply with complicated structural patterns in the data.
Data Nesting Depth Analysis
In Figure 8(a), we present the win rate for top-performing models against varying levels of maximum depth in the DAG structure, which corresponds to the longest nested data dependency flow in a sample. We observe that all models perform well for samples with maximum single nesting depth. However, the performance drops sharply with depths of two or more suggesting that long nested sequences present difficult scenarios for current models.
Total Data Dependency Analysis
In Figure 8(b), we present the win rate compared to the total number of data dependencies within a sample (a representation of the complexity of the sequence). The trends here are similar to the ones we observed in the Nesting Depth Analysis section. Models perform well for singular nested data dependency in the sample, achieving  60% Win Rate, but the performance drops sharply with two or more nested data dependencies.
Nesting Patterns Analysis
We identify common nesting patterns and analyze model performances for individual nesting patterns. Results are shown in Figure 8(c). We observe that for a simple pattern such as A $\rightarrow$ B, where the output of A is used as input by B, all models perform fairly well. However, for complex patterns such as {A, B} $\rightarrow$ C, where the outputs of both A and B are used as input by C, model performance decreases significantly. This suggests that models currently struggle with more complex patterns present in NESTful.
5 Challenges
Results show that NESTful posed a challenge for state-of-the-art LLMs for several reasons.
Data-type Adherence for the Input/Output Parameters   In the API specification, we define the data type for all parameters— input and output. The ‘type’ field specifies the data type, such as string, number, list, etc. Since APIs follow a strict structure for both input and output, it is crucial for the model to adhere to these specified format. If the model fails to follow this, especially in cases involving nested functions where the output of one API is used as the input for another, the process will fail if the output type doesn’t match the expected input type. Additionally, in nested scenarios, the output parameters of the referenced function must be used correctly; otherwise, incorrect parameter assignments will result in an incorrect answer.
Variable Assignments   As discussed in Section 3.1, we add variable assignments for each API in the output to manage parallel function calls, which is very common in real-life applications. Below is an example of parallel nested function calls:
{"input": "What is the difference between the squares of 4 and 3?"
"output": [
{"name": "square", "arguments": {arg_0: 4},
"label": "\$var_1"},
{"name": "square", "arguments": {arg_0: 3},
"label": "\$var_2"},
{"name": "subtraction", "arguments": {arg_0: \$var_1.result\$, arg_1: \$var_2.result\$},
"label": "var_3"},
In the above example, the first two functions need to be the same with different parameter values. However, for the third function subtraction, it is crucial to distinguish the outputs of the first two functions appropriately to obtain the correct result. This adds complexity to the dataset, as the models are not trained with this schema and must follow the instructions precisely. Our qualitative studies (discussed in Section 4.4) suggest the same as well.
Implicit API calling   Implicit function calling refers to a scenario where the system must invoke a specific API, along with potentially other APIs, to fulfill a user query, even though the query does not explicitly mention the task that requires that particular API. Figure 1 demonstrates an example of implicit function calling, where the user query presents an arithmetic problem without explicitly stating the APIs involved. In this case, the model must understand the problem, identify the relevant formulas, choose the appropriate function calls in sequence based on those formulas, and then populate the parameter values by either extracting information from the query or referencing the output of previous function calls. This increases the dataset complexity and makes it harder for the models.
6 Conclusion
In this work, we introduced NESTful, a new benchmark for evaluating the LLMs on nested sequences of API function calling. Existing LLMs perform poorly on this dataset as compared to their performance on existing benchmarks and identified their several modes of failure. In addition, we outlined the many challenges this dataset poses to LLM function calling approaches. By making this dataset available publicly under a permissive open-source license, we aim to push API function calling capabilities in new directions and unlock solutions to more realistic and challenging tasks.